# BarrierBench : Evaluating Large Language Models for Safety Verification in Dynamical Systems

### v1

@@ 7 Conclusion
We present a novel LLM agentic framework for barrier certificate synthesis that addresses fundamental limitations in existing approaches through natural language reasoning and adaptive mathematical structure exploration. By leveraging the reasoning capabilities of LLMs augmented with relevant barrier retrieval and verification-guided refinement, our framework automates the traditionally manual process of template selection and enables the discovery of diverse barrier certificate structures that fixed-template approaches can not discover. Moreover, we extend the framework to simultaneous barrier-controller co-synthesis, ensuring mathematical compatibility between safety certificates and feedback control laws. We introduce BarrierBench, a benchmark of 100 dynamical systems spanning continuous-time, discrete-time, linear, nonlinear, and controlled systems across various dimensions. Our evaluation demonstrates 90% success rate of the framework on this benchmark.
This work presents an initial step toward integrating language-based reasoning with formal safety verification and control synthesis. While the proposed framework and BarrierBench benchmark provide a concrete foundation for exploring this direction, we do not claim completeness or exhaustive coverage of all system classes, synthesis strategies, or verification settings. Instead, our goal is to establish a starting point and an open, extensible database that the community can collectively expand and refine. By enabling shared experimentation and collaborative benchmarking, we hope to accelerate progress toward more general, interpretable, and automated methods for safety assurance in dynamical systems. The authors remain open to any suggestions, feedback, and contribution of new systems to further develop and enrich BarrierBench.

### v2

@@ 6 Conclusion
We presented an LLM-agentic framework for barrier certificate synthesis that combines natural-language reasoning, retrieval of relevant prior barriers, and verification-guided refinement to address key limitations of existing methods. The framework automates much of the manual process of template selection, discovers diverse barrier structures beyond fixed-template approaches, and extends to barrier–controller co-synthesis for controlled systems. We also introduced BarrierBench, a benchmark of 100 dynamical systems spanning continuous-time, discrete-time, linear, nonlinear, and controlled settings, on which the framework achieves a 90% success rate.
This work takes an initial step toward integrating language-based reasoning with formal safety verification and control synthesis. Rather than claiming exhaustive coverage, we aim to provide a concrete and extensible starting point for future work. We hope that releasing both the framework and BarrierBench will support shared experimentation and accelerate progress toward more general, interpretable, and automated methods for safety assurance in dynamical systems.
Acknowledgements
This research was supported in part by the National Science Foundation under CAREER Award CCF-2146563. Ashutosh Trivedi is a Royal Society Wolfson Visiting Fellow and gratefully acknowledges support from the Wolfson Foundation and the Royal Society. The research of Sadegh Soudjani is supported by ERC 101089047 and by EIC 101070802. We also acknowledge the use of the Excalidraw tool (https://excalidraw.com) in creating the illustration in Figure 1.